# Impact of the carbon price on credit portfolio’s loss with stochastic collateral

### v1

@@ Abstract
In this work, we propose models to quantify the distortion a credit portfolio (expected and unexpected) losses, when the obligor companies as well as their guarantees belong to an economy subject to the climate transition. The economy is driven by its productivity which is a multidimensional Ornstein-Uhlenbeck process while the climate transition is declined thanks to the carbon price, a continuous deterministic process. We define each loan’s loss at default as the difference between Exposure at Default (EAD) and the liquidated collateral, which will help us to define the Loss Given Default (LGD) – the expected percentage of exposure that is lost if a debtor defaults. We consider two types of collateral: financial asset such as invoices, cash, or investments or physical asset such as real estate, business equipment, or inventory. First, if it is a financial asset, we model the later by the continuous time version of the discounted cash flows methodology, where the cash flows SDE is driven by the instantaneous output growth, the instantaneous growth of a carbon price function, and an arithmetic Brownian motion. Secondly, for physical asset, we focus on the example of a property in housing market. Therefore, we define, as Sopgoui, (2024), its value as the difference between the price of an equivalent efficient building following an exponential Ornstein-Uhlenbeck as well as the actualized renovation costs and the actualized sum of the future additional energy costs due to the inefficiency of the building, before an optimal renovation date which depends on the carbon price process. Finally, we obtain how the loss’ risk measures of a credit portfolio are skewed in the context of climate transition through carbon price and/or energy performance of buildings when both the obligors and their guarantees are affected. This work provides a methodology to calculate the (statistics of the) loss of a portfolio of secured loans, starting from a given climate transition scenario described by a carbon price.
@@
    $\mathbf{1}:=(1,\ldots,1)\in\mathbb{R}^{I}$.
@@
    $\mathbb{R}^{n\times d}$ is the set of real-valued $n\times d$ matrices ($\mathbb{R}^{n\times 1}=\mathbb{R}^{n}$), $\mathbf{I}_{n}$ is the identity $n\times n$ matrix.
@@
    $x^{i}$ denotes the $i$-th component of the vector $x\in\mathbb{R}^{d}$. For all $A:=(A^{ij})_{1\leq i,j\leq n}\in\mathbb{R}^{n\times n}$, we denote by $A^{\top}:=(A^{ji})_{1\leq i,j\leq n}\in\mathbb{R}^{n\times n}$ the transpose matrix, and $\lambda(A)$ denotes the spectrum of $A$.
@@
    $\bigotimes$ is the Kronecker product.
  • 7.
    For a given finite set $S$, we define as the cardinal of $S$, #$\#S$.
  • 8.
@@
    $(\Omega,\mathcal{H},\mathbb{P})$ is a complete probability space.
  • 10.
  • 11.
@@
    If $f:\mathbb{R}\to\mathbb{R},t\mapsto f(t)$ is a differentiable function, we note $\dot{f}$ its first derivative.
Introduction
@@
When an obligor (firm, government, or individual) defaults, the creditor (bank) stands to lose its money. One way to ensure the stability of the banks’ business and more generally the soundness of the whole financial system, is ideally, to anticipate when the default will happen and how much a bank could lose. In order to achieve that, the Basel Committee on Banking Supervision, (2017) introduces four parameters: the probability of default (PD) which measures the default risk associated with each borrower, the exposure at default (EAD) which quantifies the outstanding debt at the time of default, the loss given default (LGD) which captures the expected percentage of EAD that is lost if the debtor defaults, and the effective maturity T which represents the duration of the credit. By using, among others, these parameters, banks can compute various risk measures (such as Expected, Unexpected and Stressed Losses) which help later on to determine provisions, as well as economic and regulatory capital. An essential part of a bank risk division is to estimate how these risk measures change with various factors such as time and economic conditions.
@@
Let us focus for example on LGD. When a debtor defaults, banks can lose part or whole of its exposure. The fraction of the loss relative to EAD is LGD while the recovery rate is the fraction of EAD recovered so that $LGD=1-Recovery$. So modelling LGD or modelling recovery are equivalent. According to Chalupka and Kopecsni, (2008), there are three ways to handle LGD: ”Market LGD is observed from market prices of defaulted bonds or marketable loans soon after the actual default event. Workout LGD is derived from a set of estimated cash flows resulting from a workout and collection process, properly discounted to a date of default. Thirdly, implied market LGD is derived from risky but not defaulted bond prices using a theoretical asset pricing model”. In the IRB approach, LGD refers to Workout LGD and there are several techniques to model it. In economic modeling, as detailed by Bastos, (2010); Roncalli, (2020), LGD is a (linear or non linear) function of many factors which can be factors external to the issuer, specific to the issuer or specific to the debt issuance. That function can be obtained/calibrated through logistic regression, regression trees, or neural networks. In stochastic modeling, it is assumed that LGD follows a given distribution (parametric or non-parametric). In this case, LGD is commonly modeled by a Beta distribution as Roncalli, (2020)[Page 193] and Chalupka and Kopecsni, (2008). Its parameter are then estimated on historical data. Fermanian, (2020), for his part, proposes a joint modelling of PD and LGD by writing the potential loss at default as the difference between the debt amount (EAD) and the assets at the default date.
@@
There are secured and unsecured loans. In these approaches, even for secured loans, there is one parameter essential that is sometimes overlooked: the collateral. However, not all borrowers put up collateral when taking out loans. It is even worse, there is even some evidence that loans with collateral attached may be riskier for lenders (see Berger and Udell, (1990)). If the loan is secured, when the counterpart defaults, the bank liquidates the collateral (guarantee), and if the EAD is not reached, it can recover the remainder by liquidating other assets (called residual recovery). These guarantees can be tangible assets (buildings, business equipments, inventories, etc.) or intangible assets (cash deposits, public bonds, securities, etc.) as noted by Berger and Udell, (1990), Blazy and Weill, (2013). In the presence of collateral, the recovery (i.e. $1-\mathrm{LGD}$) is therefore made up of the value of the collateral at the date of default and the value of the residual recovery Frontczak and Rostek, (2015), and Pelizza and Schenk-Hoppé, (2020). We will model here two examples of guarantees: either a security or a (commercial or residential) building, which both, will be affected by the climate transition at its liquidation.
@@
A security can represent ownership in a corporation in the form of stock, a creditor relationship with a governmental body or a corporation represented by owning that entity’s bond; or rights to ownership as represented by an option. A security generate a stream of cash flows. The proxy of the security value is the infinite sum of the present value of the future cash flows. As we already know from the literature and from what we propose in Bouveret et al., (2023), the security (notably if it is a firm) value will be affected by transition risk. We will therefore revisit the results of Bouveret et al., (2023) where carbon price dynamics affect the firm value and credit risk measures such as probability of default, expected and unexpected losses. In particular, we redesign the multisectoral model with carbon price, the firm valuation model, and the credit risk model proposed in continuous time.
@@
In the same way, a commercial or residential building price will be affected by transition for example through the Energy Performance (or Energy Efficiency) as mentioned in Aydin et al., (2020); Franke and Nadler, (2019). Ter Steege and Vogel, (2021) quantify the depreciation by writing the price difference per square meter between two properties with different energy efficiency as the sum of the discounted value of (expected) energy cost differences. Sopgoui, (2024) enhance this work by assuming that initially, the building’s owner incurs additional energy costs (writing as a function of the carbon price) due to the inefficiency of their property. Then, he may decide to spend money on renovations to make their building energy efficient. After the renovation, they no longer incur additional energy costs. Therefore, he obtains the value of a building as the difference between the price of an equivalent efficient building following an exponential Ornstein-Uhlenbeck as well as the actualized renovation costs and the actualized sum of the future additional energy costs.
@@
The rest of the present work is organized as follows. We revisit in Section 2 the results of Bouveret et al., (2023) in a continuous time setting, namely a multisectoral economic model with carbon price, a firm valuation model, and a credit risk model. In section 3, we define the loss at default as the difference between EAD and the liquidated collateral, which will help us to define LGD. If the collateral is a financial asset, we model it in Subsection 3.2 by the continuous time version of the discounted cash flows, where the cash flows SDE is driven by the instantaneous consumption growth, the instantaneous growth of a carbon price’s function and a Brownian motion. If the collateral is a building, we will use the housing valuation under climate transition proposed in Sopgoui, (2024) to compute the portfolio’s loss. The last section is dedicated to estimations, simulations, and discussion. Our simulations show that (1) expected and unexpected losses increase when the price of carbon rises, (2) the presence of collateral significantly reduces expected and unexpected losses, (3) the positive effect of collateral on losses is reduced if the collateral is energy inefficient (for a building) or depends on a polluting sector (financial asset).
@@ 1 The problem
We consider a bank credit portfolio composed of $N\in\mathbb{N}^{*}$ firms in a closed economy (in other words no import and no export). In credit risk assessment, one of the first steps is to create homogeneous sub-portfolios of firms. As we are dealing here with climate transition risk, we would like to classify firms by carbon intensity: firms with similar carbon intensities belong to a same homogeneous sub-portfolio. It should be noted that in the absence of a climate transition, firms are traditionally clustered in terms of industry, geography, size, and credit rating, for example.
We thus assume $I\in\mathbb{N}^{*}$ ($I\leq N$) homogeneous carbon emission sectors in the economy. Nevertheless, as we rarely have the firm individual carbon emissions/intensities, we assume that each company has the carbon intensity of its industry sector. This amounts to grouping ”industry sectors” into $I$ ”carbon emission sectors”. From now on, sectors are to be interpreted as carbon emission sectors.
@@
The bank’s potential loss caused by a firm depends essentially on the default date and on the liquidation of the guarantees if they exist. The firm as well as the guarantee belong to the same economy subject to the climate transition. Thus, we build in the first stage a dynamic, stochastic, and multisectoral economic model in which direct and indirect GHG emissions from companies as well as direct GHG emissions from households are charged. We choose a representative firm in each sector and a representative household for the whole economy. By observing that each firm belongs to a sector and its cash flows are a proportion of its sales. The latter are themselves a proportion of the sectoral output, we obtain the cash flows dynamics that we use to model the value of firms in an environment where GHG emissions are charged. Then, starting from a default model in which a company defaults if its value falls below its debt, we calculate the probability of default of each firm. Finally, we compute the distortion of the (associated statistics of) loss – defined as the difference between the exposure and the liquidated collateral if exists – by the introduction of a carbon price.
@@
In this section, we revisit the framework developed in Bouveret et al., (2023) in continuous time. Precisely, we decline, in continuous time, the two standing assumptions as well as the three main results respectively on the dynamic stochastic multisectoral model with carbon emissions costs, on the firm valuation model, and on the structural credit risk model. Most of the proofs can be derived from the discrete time so that we will either skip them or detail them in appendix.
@@ 2.1 A Multisectoral Model with Carbon price
Each sector $i\in\mathcal{I}$ has a representative firm which produces a single good, so that we can associate sector, firm and good. We introduce the following standing assumption which describes the productivity, which is considered to have stationary Ornstein-Uhlenbeck dynamics.
@@
where $(B_{t}^{\mathcal{Z}})_{t\in\mathbb{R}^{*}}$ is a $I$-dimensional Brownian Motion, and where the constants $\mu,\mathcal{A}_{0}\in\mathbb{R}^{I}$, the matrices $\Gamma,\Sigma\in\mathbb{R}^{I\times I}$, $\mathcal{Z}_{0}\sim\mathcal{N}\left(0,\Sigma\Sigma^{\top}\right)$, and $0<\varsigma\leq 1$ is an intensity of noise parameter that is fixed: it will be used later to obtain a tractable proxy of the firm value. Moreover, $\Sigma$ is a positive definite matrix and $-\Gamma$ is a Hurwitz matrix i.e. its eigenvalues have strictly negative real parts.
The processes $\mathcal{Z}^{i}$ and $\mathcal{A}^{i}$ play a major role in our factor productivity model since, for any $i\in\mathcal{I}$, the total factor productivity of sector $i$ is defined as
$A^{i}:=\exp{(\mathcal{A}^{i})},$ (2.2)
so that $\mathcal{Z}^{i}$ is the log-productivity growth and $\mathcal{A}^{i}$ is the cumulative log-productivity growth. In the rest of the paper, the terminology ”productivity” will be used within a context that will allow the reader to understand if the term refers to $\mathcal{Z}^{i}$, $\mathcal{A}^{i}$, or $A^{i}$.
We also introduce the following filtration $\mathbb{G}:=(\mathcal{G}_{t})_{t\in\mathbb{R}^{*}}$ with $\mathcal{G}_{0}:=\sigma(\mathcal{Z}_{0})$ and for $t>0$, $\mathcal{G}_{t}:=\sigma\left(\left\{\mathcal{Z}_{0},B_{s}^{\mathcal{Z}}:s\leq t\right\}\right)$.
Remark 2.2 (O.U. process).
We have the following results on O.U. that we will use later:
@@
    Since $-\Gamma$ is a Hurwitz matrix, then if we note $\lambda_{\Gamma}:=\max_{\lambda\in\lambda(\Gamma)}Re(\lambda)$, there exists $c_{\Gamma}>0$ so that $\lVert e^{-\Gamma t}\rVert<c_{\Gamma}e^{-\lambda_{\Gamma}t}$ for all $t\geq 0$. Therefore, according to Gobet and She, (2016)[Proposition 2], $\mathcal{Z}$ has a unique stationary distribution which is Gaussian with mean 00 and covariance $\int_{0}^{+\infty}e^{-\Gamma u}\Sigma\Sigma^{\top}e^{-\Gamma^{\top}u}\mathrm{d}u$.
@@
    and conditionally on $\mathcal{G}_{t}$, $\mathcal{A}_{t+h}$ has an $I$-dimensional normal distribution with the mean vector
    $M^{\mathcal{A},h}_{t}:=\mu h+\varsigma\Upsilon_{h}\mathcal{Z}_{t}+\mathcal{A}_{t},$ (2.5)
    with
@@
    For later use, we define
    $\mathcal{A}^{\circ}_{t}:=\mathcal{A}_{t}-\mathcal{A}_{0},$ (2.8)
    and observe that $(\mathcal{A}^{\circ}_{t},\mathcal{Z}_{t})_{t\geq 0}$ is a Markov process.
Firms emit GHG when they consume intermediary input from other sectors and when they produce output. Likewise, households emit GHG when they consume. All these emissions are charged through a carbon price dynamics. For the whole economy, we introduce a deterministic and exogenous carbon price in euro/dollar per ton. It allows us to model the impact of the transition pathways on the whole economy. We will note $\delta$ the complete carbon price process. We shall then assume the following setting.
@@
    Let $0\leq t_{\circ}<t_{\star}$ be given. The sequence $\delta$ satisfies
@@
      for $t\in[0;t_{\circ}]$, $\delta_{t}=\delta_{0}\in(\mathbb{R}_{+})^{I}$, namely the carbon price is constant;
@@
      for $t\in(t_{\circ},t_{\star})$, $\delta_{t}\in(\mathbb{R}_{+})^{I}$, the carbon price may evolve;
@@
      for $t\geq t_{\star}$, $\delta_{t}=\delta_{t_{\star}}\in(\mathbb{R}_{+})^{I}$, namely the carbon price is constant.
    We assume moreover that $t\mapsto\delta_{t}$ is $\mathcal{C}^{1}(\mathbb{R}_{+},\mathbb{R}_{+})$.
@@
$\mathfrak{d}_{t}:=(\tau_{t}\delta_{t},\zeta_{t}\delta_{t},\kappa_{t}\delta_{t}).$ (2.11)

### v2

@@ Abstract
The aim of this work is to propose an end-by-end modeling framework to evaluate the risk measures of a bank’s portfolio of collateralized loans when the bank, the borrowers, as well as their guarantees operate in an economy subject to the climate transition. The economy, organized in sectors, is driven by its productivity which is a multidimensional Ornstein-Uhlenbeck process while the climate transition is declined thanks to the carbon price and carbon intensities, continuous deterministic processes. We thus derive the dynamics of macroeconomic variables for each climate transition scenario. By considering that a firm defaults if it is over-indebted i.e. its market value – depreciated due to the carbon price – becomes less than the market value of its debt, we define each loan’s loss at default as the difference between Exposure at Default (EAD) and the liquidated collateral, which will help us to define the Loss Given Default (LGD) – the expected percentage of exposure that is lost if a debtor defaults. We consider two types of collateral: financial asset such as invoices, cash, or investments or physical asset such as real estate, business equipment, or inventory. First, if it is a financial asset, we model the later by the continuous time version of the discounted cash flows methodology, where the cash flows growth is driven by the instantaneous output growth, the instantaneous growth of a carbon price function, and an arithmetic Brownian motion. Secondly, for physical asset, we focus on the example of a property in housing market. Therefore, we define, as Sopgoui, (2024), its value as the difference between the price of an equivalent efficient building following an exponential Ornstein-Uhlenbeck (OU) as well as the actualized renovation costs and the actualized sum of the future additional energy costs due to the inefficiency of the building, before an optimal renovation date which depends on the carbon price process. Finally, we obtain expressions for risk measures of a portfolio of collateralized loans as a function of various parameters and variables, mainly those linked to the climate transition, such as the carbon price and the energy efficiency of buildings. These risk measures will be used by banks, depending on the climate transition scenarios, to define their operating expenses, the fees applied to clients, as well as their economic and regulatory capital.
@@
    For $n,d\in\mathbb{N}^{*}$, $\mathbb{R}^{n\times d}$ is the set of real-valued $n\times d$ matrices ($\mathbb{R}^{n\times 1}=\mathbb{R}^{n}$), $\mathbf{I}_{n}$ is the identity $n\times n$ matrix and $\mathbf{1}:=(1,\ldots,1)\in\mathbb{R}^{n}$.
@@
    $x^{i}$ denotes the $i$-th component of the vector $x\in\mathbb{R}^{d}$. For $A:=(A^{ij})_{1\leq i,j\leq n}\in\mathbb{R}^{n\times n}$, we denote by $A^{\top}:=(A^{ji})_{1\leq i,j\leq n}\in\mathbb{R}^{n\times n}$ the transpose matrix, and $\lambda(A)$ denotes the spectrum of $A$.
@@
    For a given finite set $S$, we define as the cardinal of $S$, #$\#S$.
@@
  • 7.
    $(\Omega,\mathcal{H},\mathbb{P})$ is a complete probability space.
  • 8.
@@
  • 10.
  • 11.
    If X is a log-Normal random variable, we note $X\sim\mathcal{LN}(\mu,\sigma^{2})$ with $\mu:=\mathbb{E}[\log(X)]$ (the mean of $\log(X)$) and $\sigma^{2}:=\mathbb{V}[\log(X)]$ (the variance of $\log(X)$).
@@
    For $A\subset\mathbb{R}$ and $B\subset\mathbb{R}^{I}$, if $f:A\to B,t\mapsto f(t)$ is a differentiable function, we note $\dot{f}$ its first derivative. We also note $\mathfrak{D}(A,B)$ the set of derivative functions from $A$ to $B$.
1 Introduction
There is now evidence that climate change has adverse effects on the environment, on global warming, and on human societies. Climate risk has two components: physical risk which arises from alterations in climate variables, like increasing temperatures and melting ice, or from extreme weather events, such as droughts or typhoons that can damage infrastructure and endanger populations; and transition risk which comes from the necessity of transitioning to a low-carbon economy, leading to the implementation of regulatory/political measures, potential technological changes, and the evolution of consumer preferences. We focus in this work on transition risk and particularly on political changes. The most well-known way to that is the carbon price. This means that the greenhouse gases (GHG) emissions of economic agents are charged. For instance, the carbon price can modify the three main components of credit risk, which are the borrower’s cash flows, the value of his assets, and the value of his guarantees. The aim of this work is to compute the losses of a credit portfolio when the bank, its borrowers, and their guarantees operate in an economy that undergoes the climate transition.
@@ 1 Introduction
When an obligor (firm, government, or individual) defaults, the creditor (bank) stands to lose its money. One way to ensure the stability of the banks’ business and more generally the soundness of the whole financial system, is ideally, to anticipate when the default will happen and the amount a bank could lose. In order to achieve that, the Basel Committee on Banking Supervision, (2017) introduces four parameters: the probability of default (PD) which measures the default risk associated with each borrower, the exposure at default (EAD) which quantifies the outstanding debt at the time of default, the loss given default (LGD) which captures the expected percentage of EAD that is lost if the debtor defaults, and the effective maturity T which represents the duration of the credit. By using these parameters, banks can compute various risk measures (such as expected, unexpected, and stressed losses) which help later on to determine provisions, as well as economic and regulatory capital. Using Gordy, (2003), we can obtain the losses as a function of PD, LGD, EAD, and T. An essential part of a bank risk division is to estimate how the risk measures change with various factors such as time and economic conditions.
@@
In Bouveret et al., (2025), we propose a methodology in discrete time without collateral (i.e., with LGD independent of the economy and climate transition). However, when a debtor defaults, the bank can lose all or part of its exposure. The fraction of the loss relative to EAD is LGD while the recovery rate is the fraction of EAD recovered so that $LGD=1-Recovery$. So modeling LGD or recovery is equivalent. According to Chalupka and Kopecsni, (2008), there are three types of LGD: ”market LGD is observed from market prices of defaulted bonds or marketable loans soon after the actual default event. Workout LGD is derived from a set of estimated cash flows resulting from a workout and collection process, properly discounted to a date of default. Thirdly, implied market LGD is derived from risky but not defaulted bond prices using a theoretical asset pricing model”. When consider the IRB approach [^1] Internal Rating Based (see Basel Committee on Banking Supervision, (2017)), LGD refers to Workout LGD and there are several techniques to model it. In economic modeling, as detailed by Bastos, (2010); Roncalli, (2020), LGD is a function of different factors which can be external to the issuer, specific to the issuer, or specific to the debt issuance. That function can be obtained through logistic regression, regression trees, or neural networks. In stochastic modeling, it is assumed that LGD follows a given (parametric or non-parametric) distribution. For example, LGD is commonly modeled by a Beta distribution as Roncalli, (2020)[Page 193] and Chalupka and Kopecsni, (2008).Fermanian, (2020), for his part, proposes a joint modeling of PD and LGD by writing the potential loss at default as the difference between the debt amount (EAD) and the assets at the default date.
@@
In a credit portfolio, there are secured (or collateralized ) and unsecured loans. We call secured loans when there is collateral (also called guarantee) and unsecured loans when there is not. Of course, not all borrowers put up collateral when taking out loans, and there is evidence that loans with collateral may be riskier for lenders. Berger and Udell, (1990) note that banks require more collateral from riskier borrowers, and although seizing a guarantee when a borrower defaults reduces the bank’s loss, this does not always fully compensate the bank for the fact that the loan was initially riskier. For a secured loan, when the counterpart defaults, the bank liquidates the collateral, and if the EAD is not reached, it can recover the remaining amount by liquidating other assets (called residual recovery). These guarantees can be tangible assets (buildings, business equipments, inventories, etc.) or intangible assets (cash deposits, public bonds, securities, etc.) as noted by Berger and Udell, (1990), Blazy and Weill, (2013). For secured loans, the recovery (that is, $1-\mathrm{LGD}$) is therefore made up of both the value of the collateral at the date of default and the value of the residual recovery Frontczak and Rostek, (2015), and Pelizza and Schenk-Hoppé, (2020). However, in most LGD modeling approaches, the presence or absence of the guarantees is often overlooked. We will tackle these limits here by considering two examples of guarantees: either a security or a (commercial or residential) building, which both will be affected by the climate transition.
@@
A security can represent ownership in a corporation in the form of stock, a creditor relationship with a governmental body, or a corporation represented by owning that entity’s bond; or rights to ownership as represented by an option. A security generates a stream of cash flows. The proxy of the security value is the (in)finite sum of the present value of the future cash flows. As we already know from what we propose in Bouveret et al., (2025), the value of the security (particularly if it is a firm) will be affected by the transition risk. We will therefore revisit the results of Bouveret et al., (2025) where carbon price dynamics affects the firm value and credit risk measures such as probability of default, expected and unexpected losses. In particular, we redesign the multisectoral model with carbon price, the firm valuation model, and the credit risk model proposed in continuous time.
@@
In the same way, a commercial or residential building price will be affected by the climate transition, for example through Energy Performance (or Energy Efficiency) as mentioned in Aydin et al., (2020); Franke and Nadler, (2019). Ter Steege and Vogel, (2021) quantify the depreciation by writing the price difference per square meter between two properties with different energy efficiency as the sum of the discounted value of the (expected) energy cost differences. Sopgoui, (2024) enhances this work by assuming that initially the owner of the building incurs additional energy costs (written as a function of the carbon price) due to the inefficiency of his property. Then, he may decide to spend money on renovations to make his building energy efficient. After the renovation, he no longer incurs additional energy costs. Therefore, he obtains the value of his building as the difference between the price of an equivalent efficient building following an exponential Ornstein-Uhlenbeck as well as the actualized renovation costs and the actualized sum of the future additional energy costs.
Thanks to our micro-founded approach to a credit portfolio, we obtain explicit expressions of risk measures (especially probability of default, loss given default, expected and unexpected losses) depending on variables and parameters traditionally found in the literature, such as the the borrower’s cash flows, its exposure, the calculation time, the nature of the collateral, the liquidation time, date, and costs, the interest rate, the productivity, etc. Additionally, they now depend on climate transition-specific parameters, such as carbon price, carbon intensities, energy efficiency, or renovation costs. In summary, our framework enables the calculation of the potential loss of a collateralized loan and also measures the variation in this loss depending on the borrower’s and the collateral’s virtuousness in terms of GHG emissions. Precisely, we will show that (1) expected and unexpected losses increase when the price of carbon increases, (2) the presence of collateral significantly reduces expected and unexpected losses, and (3) the positive effect of collateral on losses is reduced if the collateral is energy inefficient (for a building) or depends on a polluting sector (financial asset). And at each stage, variations can be precisely quantified.
@@
The rest of the present work is organized as follows. In section 2, we clarify the problem that this paper aims to solve. We revisit in  section 3 the results of Bouveret et al., (2025) in a continuous time setting, namely a multisectoral economic model with carbon price, a firm valuation model, and a credit risk model. In section 4, we define the loss at default as the difference between EAD and the liquidated collateral, which will help us to define LGD. If the collateral is a financial asset, we model it in section 4.2 by the continuous time version of the discounted cash flows, where the cash flows growth is driven by the instantaneous consumption growth, the instantaneous growth of a carbon price’s function and a Brownian motion. If the collateral is a building, we will use the housing valuation under the climate transition proposed in Sopgoui, (2024) to compute the loss of the portfolio. The last sections (5 and 6) are dedicated to estimations, simulations, and discussion.
@@ 2 The problem
We consider a bank credit portfolio composed of $N\in\mathbb{N}^{*}$ firms in a closed economy (in other words, no import and no export). In credit risk assessment, one of the first steps is to create homogeneous sub-portfolios of firms. As we are dealing here with climate transition risk, we would like to classify firms by carbon intensity – the quantity of GHG in tons emits for each unit of production (in ton per euro) [^2] To obtain this, one could simply divide each company’s greenhouse gas emissions by its revenue. – so that firms with similar carbon intensities belong to a same homogeneous sub-portfolio. It should be noted that in the absence of a climate transition, firms are traditionally clustered in terms of industry, geography, size, and credit rating, for example.
We thus assume $I\in\mathbb{N}^{*}$ ($I\leq N$) homogeneous carbon emission sectors in the economy. Nevertheless, as we rarely have the firm individual carbon emissions/intensities (especially for small and medium enterprises), we assume that each company has the carbon intensity of its industry sector. This amounts to grouping ”industry sectors” into $I$ ”carbon intensity sectors”. From now on, sectors are to be interpreted as carbon emission sectors.
@@
The bank’s potential loss caused by a firm depends essentially on the default date and on the liquidation of the guarantees if they exist. The firm as well as the guarantee belong to the same economy which is subject to the climate transition. Thus, we build in the first stage a dynamic, stochastic, and multisectoral economic model in which direct and indirect GHG emissions from companies as well as direct GHG emissions from households are charged. We choose a representative firm in each sector and a representative household for the whole economy. By observing that each firm belongs to a sector and its cash flows are a proportion of its sales. The latter are themselves a proportion of the sectoral output. We obtain the cash flows dynamics that we use to determine the value of firms in an environment where GHG emissions are charged. Then, starting from a default model in which a company defaults if its value falls below its debt, we calculate the probability of default of each firm. Finally, considering that the collateral is a GHG-emitting asset in the economy, we compute the distortion of the (associated statistics of) loss – defined as the difference between the exposure and the liquidated collateral – by the introduction of a carbon price.
@@
In Bouveret et al., (2025), we propose an end-to-end methodology, starting from a transition scenario modeled by the carbon price and carbon intensities, to the impact on different credit risk measures – especially the probability of default, the expected and unexpected losses – with both EAD and LGD deterministic and without collateral. In this section, we revisit that framework in continuous time. Precisely, we decline, in continuous time, the two standing assumptions as well as the three main results respectively on the dynamic stochastic multisectoral model with carbon emissions costs, on the firm valuation model, and on the structural credit risk model. Most of the proofs can be derived from the discrete time so we will skip them or detail them in appendix B.
@@ 3.1 A Multisectoral Model with Carbon price
The goal here is to derive the dynamics of macroeconomic variables (GHG emissions, output, consumption, labor, and intermediary inputs) per sector and per climate transition scenario.
Each sector $i\in\mathcal{I}$ has a representative firm which produces a single good, so that we can associate sector, firm, and good. We introduce the following standing assumption which describes the productivity, which is considered to have stationary Ornstein-Uhlenbeck dynamics.
@@
where $(B_{t}^{\mathcal{Z}})_{t\in\mathbb{R}_{+}}$ is a $I$-dimensional standard Brownian Motion, and where the constants $\mu,\mathcal{A}_{0}\in\mathbb{R}^{I}$, the matrices $\Gamma,\Sigma\in\mathbb{R}^{I\times I}$, $\mathcal{Z}_{0}\sim\mathcal{N}\left(0,\Sigma\Sigma^{\top}\right)$, and $0<\varsigma\leq 1$ is an intensity of noise parameter that is fixed: it will be used later to obtain a tractable proxy of the firm value. Moreover, $\Sigma$ is a positive definite matrix and $-\Gamma$ is a Hurwitz matrix i.e. its eigenvalues have strictly negative real parts.
The processes $\mathcal{Z}^{i}$ and $\mathcal{A}^{i}$ play a major role in our factor productivity model since, for any $i\in\mathcal{I}$, the total factor productivity of sector $i$ is defined as $A^{i}:=\exp{(\mathcal{A}^{i})}$, so that $\mathcal{Z}^{i}$ is the log-productivity growth and $\mathcal{A}^{i}$ is the cumulative log-productivity growth. In the rest of the paper, the terminology ”productivity” will be used within a context that will allow the reader to understand if the term refers to $\mathcal{Z}^{i}$, $\mathcal{A}^{i}$, or $A^{i}$.
We also introduce the following filtration $\mathbb{G}:=(\mathcal{G}_{t})_{t\in\mathbb{R}_{+}}$ with $\mathcal{G}_{0}:=\sigma(\mathcal{Z}_{0})$ and for $t>0$, $\mathcal{G}_{t}:=\sigma\left(\left\{\mathcal{Z}_{0},B_{s}^{\mathcal{Z}}:s\leq t\right\}\right)$.
Remark 3.2 (OU process).
We have the following results on OU that we will use later:
@@
    Since $-\Gamma$ is a Hurwitz matrix, then if we note $\lambda_{\Gamma}:=\max_{\lambda\in\lambda(\Gamma)}Re(\lambda)$, there exists $c_{\Gamma}>0$ so that $|e^{-\Gamma t}|<c_{\Gamma}e^{-\lambda_{\Gamma}t}$ for all $t\geq 0$. Therefore, according to Gobet and She, (2016)[Proposition 2], $\mathcal{Z}$ has a unique stationary distribution which is Gaussian with mean 00 and covariance $\int_{0}^{+\infty}e^{-\Gamma u}\Sigma\Sigma^{\top}e^{-\Gamma^{\top}u}\mathrm{d}u$.
@@
    and conditionally on $\mathcal{G}_{t}$, $\mathcal{A}_{t+h}$ has an $I$-dimensional normal distribution with the mean vector $M^{\mathcal{A},h}_{t}:=\mu h+\varsigma\Upsilon_{h}\mathcal{Z}_{t}+\mathcal{A}_{t}$, with
@@
    For later use, we define $\mathcal{A}^{\circ}_{t}:=\mathcal{A}_{t}-\mathcal{A}_{0}$, and observe that $(\mathcal{A}^{\circ}_{t},\mathcal{Z}_{t})_{t\geq 0}$ is a Markov process.
Firms emit GHG when they consume intermediary input from other sectors and when they produce output. Likewise, households emit GHG when they consume. All these emissions are charged through a carbon price dynamics. For the whole economy, we introduce a deterministic and exogenous carbon price in euro per ton of CO2-equivalent. It allows us to model the impact of the transition pathways on the whole economy. We shall then assume the following setting.
@@
    Let $0\leq t_{\circ}<t_{\star}$ be given. The complete carbon price sequence noted $\delta$ satisfies
@@
      for $t\in[0;t_{\circ}]$, $\delta_{t}=\delta_{0}\in\mathbb{R}_{+}$, namely the carbon price is constant;
@@
      for $t\in(t_{\circ},t_{\star})$, $\delta_{t}\in\mathbb{R}_{+}$, the carbon price may evolve;
@@
      for $t\geq t_{\star}$, $\delta_{t}=\delta_{t_{\star}}\in\mathbb{R}_{+}$, namely the carbon price is constant.
    We assume moreover that $t\mapsto\delta_{t}$ is $\mathfrak{D}(\mathbb{R}_{+},\mathbb{R}_{+})$.
@@
Remark 3.4.
At each time $t\geq 0$ and for $i\in\mathcal{I}$,
  • 1.
    $\tau_{t}^{i}$ represents the carbon intensity of the firms production of sector $i$ at time $t$;
  • 2.
    $\kappa_{t}^{i}$, the carbon intensity of the household final consumption in sector $i$ at time $t$,
  • 3.
    and for $j\in\mathcal{I}$, $\zeta^{ji}_{t}$, the carbon intensity of the firms of sector $i$ intermediate consumption in sector $j$ at time $t$.
@@
$\mathfrak{d}_{t}:=(\delta_{t}\tau_{t},\delta_{t}\zeta_{t},\delta_{t}\kappa_{t}).$ (3.8)